Problem 7-4 (Part Level Submission) From inception of operations to December 31, 2017, Skysong Corporation provided for uncollectible accounts receivable under the allowance method. The provisions are recorded, based on analyses of customers with different risk characteristics. Bad debts written off were charged to the allowance account; recoveries of bad debts previously written off were credited to the allowance account, and no year-end adjustments to the allowance account were made. Skysong’s usual credit terms are net 30 days. The balance in Allowance for Doubtful Accounts was $145,500 at January 1, 2017. During 2017, credit sales totaled $9,135,400, the provision for doubtful accounts was determined to be $91,354, $91,354 of bad debts were written off, and recoveries of accounts previously written off amounted to $17,970. Skysong installed a computer system in November 2017, and an aging of accounts receivable was prepared for the first time as of December 31, 2017. A summary of the aging is as follows. Classification by Month of Sale Balance in Each Category Estimated % Uncollectible November–December 2017 $1,184,700 2% July–October 696,600 10% January–June 435,900 23% Prior to 1/1/17 161,800 75% $2,479,000 Based on the review of collectibility of the account balances in the “prior to 1/1/17” aging category, additional receivables totaling $65,900 were written off as of December 31, 2017. The 75% uncollectible estimate applies to the remaining $95,900 in the category. Effective with the year ended December 31, 2017, Skysong adopted a different method for estimating the allowance for doubtful accounts at the amount indicated by the year-end aging analysis of accounts receivable.

Collapse question part

(a)

Prepare a schedule analyzing the changes in Allowance for Doubtful Accounts for the year ended December 31, 2017. Show supporting computations in good form. (Hint: In computing the 12/31/17 allowance, subtract the $65,900 write-off.)

These are the accounts to be used:

Provisions for doubtful accounts

Balance at December 31, 2017 adjusted

Recovery in 2017 of bad debts written off previously

Balance at January 1, 2017

Write-offs for 2017

Balance at December 31, 2017 before change in accounting estimate

Increase due to change in accounting estimate during 2017
Answer #1

Answer:

Schedule analyzing the changes in Allowance for Doubtful Accounts for the year ended December 31, 2017 is as follows:

Skysong Schedule analyzing the changes in Allowance for Doubtful Accounts For the year ended December 31, 2017 Item Amount Ba

Supporting calculations:

Write offs during the year = $91,354 + $65,900 = $157,254

Balance at December 31, 2017 before change in accounting estimate = 145500 + 91354 - 157254 + 17970 = $97,570

Provision required based on aging of accounts receivable = Balance at December 31, 2017 adjusted

= 1184700 * 2% + 696600 * 10% + 435900 * 23% + 95900 * 75%

= $265,536

Increase due to change in accounting estimate during 2017 = Balance at December 31, 2017 adjusted - Balance at December 31, 2017 before change in accounting estimate

= 265536 - 97570

= $167,966